OpenStudy (vortish):

solve the linear inequality. 5 -7/9x > 7 - 5/1 -7/9x >= 7 -45/7x >= 7

OpenStudy (anonymous):

is that a subtraction of two inequalities?

OpenStudy (anonymous):

what is that " - " on the 2nd line?

OpenStudy (vortish):

it is suppose to be ge

OpenStudy (anonymous):

ok, no problem...

OpenStudy (vortish):

\[\ge \]

OpenStudy (anonymous):

so it's like: 5 - (7/9)x >= 7 ?

OpenStudy (vortish):

im trying to do my own work just have some one look at it to catch my mistakes

OpenStudy (anonymous):

I know... that is definitely the way to go. Step 1: move the 5 from left to right by subtracting 5 from both sides... 5 - (7/9)x >= 7 You can do it...

OpenStudy (vortish):

its \[5 -\frac{ 7 }{ 9 } x \ge 7\]

OpenStudy (anonymous):

works for me :) My step 1 above is still a good start... get the numbers on the right side by subtracting 5 from both sides.

OpenStudy (vortish):

-7/9x\[\ge 2\]

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

yeh, I got it.... -(7/9)x >= 2 Next step... solve for x by multiplying both sides by -(9/7) so you get just "x" on the left. Here's where you have to reverse the inequality.

OpenStudy (vortish):

x = -14/9

OpenStudy (vortish):

or x - 7/3

OpenStudy (anonymous):

I think you made a math error... Here's where we were... -(7/9)x >= 2 -(7/9)(-9/7)x <= 2(-9/7) multiplying both sides by -9/7 & reversing inequality x <= -18/7 Done

OpenStudy (vortish):

i see were my error is ty